Want to make 30 bounds of a bird seed that costs $3 per pound. They are using thistle seed for $2 per pound and sunflower seeds for $3.50 per pound. How many pounds of sunflower seeds should be used in this mixture?
algebra 1

Answer by checkley71(8403)

2X+3.5(30-X)=3*30
2X+105-3.5X=90
-1.5X=90-105
-1.5X=-15
X=-15/-1.5
X=10 POUNDS OF $2 SEEDS.
30-10=20 POUNDS OF $3.50 SEEDS.
PROOF
2*10+3.5*20=90
20+70=90
90=90